OpenStudy (anonymous):

f(x)=3x^2 and g(x)=-2x-5 what is the value of f(g(-4))

OpenStudy (dumbcow):

g(-4) = -2(-4)-5 = 8-5 = 3 f(3) = 3*9 = 27

OpenStudy (anonymous):

Thanks. Appreciate our help :)

OpenStudy (rajee_sam):

they are asking for the value of f (g(-4)) You have to find what f (g(x)) is first f(g(x)) = 3 ( -2x -5)² = 3 ( 4x² + 20x + 25) now f(g(-4)) = 3 ( (-4)² + (20 * -4) + 25) = 3 ( 16 - 80 + 25) = 3 (-39) = -117